10. Factorise each of the following:
(i) 27y3 + 125e' (ii) 64m8 - 343n9•

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) Consider 27y3 + 125z3 c: (3y)3 + (5z)s = cay + 5z) ((3y'f/- - C3yX5z> + C5z>2J [Using identity:.r3 + y8 • (.r + y Xr - xy + y2)J = (3y + mX9y2 - 15yz + (u) Conside.r 64m8 - 34&3 = (4m)3 - (7n"J = (4m - 7n) {(4m)2 + (4mX7n) + (7n>21 [Using identity:.r8 - y3 = (.r - y"h;2 + xy + y2)] = (4m - 7nX16m2 + 28mn + 49n2>.